# BOTS: A Unified Framework for Bayesian Online Task Selection in LLM Reinforcement Finetuning

### v2

@@ F.1 Computational Overhead
F.2 Interpolation-Based Implicit Evidence: Empirical Results
@@
To empirically validate our interpolation-based implicit evidence estimator, we assess its predictive quality against the evolving empirical success probabilities of the training model. We examine the trajectory under vanilla training, i.e., uniformly sampling tasks for training. At each step, we compare the predictions from our interpolation-based estimator with the ground-truth online task success probabilities. Two metrics are used: (i) Pearson Correlation, measuring the linear relationship between estimated and empirical difficulties, and (ii) ROC AUC, evaluating the ability to distinguish effective tasks (success strictly between 0 and 1) from ineffective tasks (success equal to 0 or 1). Evaluations are conducted throughout training for models of different scales.
@@
Figure 5: Performance of the implicit evidence estimator during training of Qwen2.5-1.5B-Instruct (Left) and Qwen2.5-7B (Right), measured by Pearson Correlation and ROC AUC.
@@
As shown in Figure 5, the estimator exhibits distinct behaviors across model scales. For Qwen2.5-1.5B-Instruct (Left), both Pearson Correlation and ROC AUC remain stable throughout training; the consistently positive correlation and ROC AUC above 0.5 demonstrate that interpolation-based implicit evidence effectively captures task difficulty. Notably, this model is weaker than both reference models (Qwen2.5-7B-Instruct and Qwen3-32B), confirming that our estimator possesses extrapolation capability. For the larger Qwen2.5-7B (Right), both Pearson Correlation and ROC AUC decline in later stages: Pearson Correlation remains positive, but ROC AUC approaches 0.5, indicating that implicit evidence becomes less informative for difficulty prediction as the model matures. This highlights the necessity of incorporating direct evaluations as explicit evidence to maintain accurate estimation.
F.3 Sampling from Posterior
@@
F.4 Dynamics of Selected Tasks
@@
F.5 Comprehensive Results on Math
Results on Math-1.5B
The results for Qwen2.5-1.5B-Instruct on the GURU-Math dataset are presented in Figure 9, Figure 10, and Table 8. The effectiveness of task selection is first evaluated through the Effective Task Ratio (ETR). BOTS demonstrates a rapid increase in ETR, peaking above 0.8 and sustaining a high level throughout training. In contrast, the random and BOTS-MoPPS baselines show minimal improvement, with ETR remaining below 0.4. While BOTS-DOTS initially performs well, it exhibits a declining ETR in later stages.
This superior task selection efficiency translates into significant gains in downstream performance. The aggregated performance plot shows that BOTS achieves a steeper learning curve and consistently outperforms the baselines. The quantitative results in Table 8 corroborate these findings. Specifically, for the aggregated performance, BOTS achieves a final TTB of 0.64, indicating a 36% reduction in training steps to reach the baseline’s best performance. It also attains a final BSF of 1.05, representing a 5% improvement in peak performance under the same training budget.
Figure 9: Qwen2.5-1.5B-Instruct on Math. Ratio of sampled training tasks with different passing rates.
Figure 10: Qwen2.5-1.5B-Instruct on Math. Performance on downstream code benchmarks (MATH500, AIME24) and their aggregation.
Benchmark MATH500 AIME24 Math Aggregated Performance
Offline 0.96 0.49 0.72 1.15 1.11 1.02 0.54 - - 1.00 2.00 0.75 0.96 0.51 0.78 1.16 1.12 1.01
BOTS-MoPPS 1.34 1.34 - 0.91 0.90 0.98 0.78 - - 1.25 1.25 0.67 1.78 1.32 - 0.90 0.90 0.99
BOTS-DOTS 0.78 0.51 0.72 1.12 1.10 1.04 0.71 1.12 1.12 0.75 1.50 1.08 0.83 0.75 0.81 1.11 1.09 1.08
BOTS 0.89 0.49 0.57 1.13 1.05 1.05 0.51 1.00 1.00 1.25 1.75 1.00 0.89 0.56 0.64 1.12 1.07 1.05
Results on Math-7B
We present the results for the Qwen2.5-7B model on the GURU-Math dataset in Figure 11, Figure 12, and Table 9. Consistent with the 1.5B experiments, our recommended BOTS setting demonstrates robust task selection efficiency. Its ETR rises rapidly to over 0.9 and is maintained at a high level, significantly outperforming the random, offline, and explicit-only baselines. This efficiency translates to tangible performance gains, with Table 9 indicating a final aggregated TTB of 0.63, confirming significant training acceleration.
A notable difference in the 7B setting is the highly competitive performance of BOT-DOTS. Unlike in the 1.5B experiments where its ETR degraded over time, this strategy maintains a high ETR throughout the training process, comparable to that of BOTS. Consequently, its downstream performance is nearly identical to our full BOTS framework, as evidenced by the overlapping performance curves and similar metrics in the corresponding figures and table. This suggests that for a highly capable base model in the math domain, the guidance from implicit evidence alone can be exceptionally effective.
@@ F.5 Comprehensive Results on Math
F.6 Comprehensive Results on Code
Results on Code-1.5B
To assess the generalizability of our framework, we replicate the comparative experiments on the GURU-Code dataset. The results, presented in Figure 13, 14 and Table 10, largely affirm the conclusions drawn from the Math domain. Our BOTS again achieves a consistently high ETR, significantly outperforming the random and offline baselines. This demonstrates its robust capability to construct an efficient curriculum across diverse domains. Correspondingly, BOTS delivers strong downstream performance, particularly in the TTB metric, indicating effective training acceleration.
However, the results on Code also reveal distinct characteristics. Most notably, BOTS-DOTS exhibits exceptionally competitive performance. As detailed in Table 10, it matches or even surpasses our recommended BOTS setting on several benchmarks, especially in the BSF metric. This suggests that the implicit evidence from our interpolator is particularly potent for the Code domain. These findings highlight that while the optimal balance of evidence sources may be domain-dependent, BOTS provides a robust and unified solution that remains a top-performing method across both settings.
@@ F.6 Comprehensive Results on Code
Results on Code-7B
We extended our analysis to the Qwen2.5-7B model to test scalability. The results, presented in Figure 15, 16 and Table 11, primarily demonstrate the consistent robustness of BOTS in optimizing the task selection. As shown in the Valid Ratio plot, our recommended setting achieves the highest proportion of valid tasks, particularly during the early and middle training stages. This consistent efficiency in task selection across different model scales confirms the robustness of our evidence-fusion approach.
On downstream benchmarks, the performance differences between methods are less pronounced, with all strategies converging to a similarly high level of performance. This outcome is expected, as the strong intrinsic capabilities of the 7B base model can diminish the performance differentiation among curriculum strategies on standard benchmarks. Nevertheless, the clear advantage of BOTS in task selection efficiency underscores its value in optimizing computational resources, even when final performance margins are narrow.
@@
F.7 Comprehensive Results on Logic
Results on Logic-1.5B
We evaluate the framework’s performance on the GURU-Logic dataset, where our recommended BOTS setting demonstrates competitive performance comparable to its results in the Math and Code domains. The curriculum dynamics in Logic, however, exhibit a unique profile. As shown in Figure 17, the Effective Task Ratio (ETR) for most strategies follows a distinct rise-and-fall pattern.
This task selection behavior is reflected in the downstream performance, detailed in Figure 18 and Table 12. The explicit-only strategy again yields highly competitive results. Against this backdrop, our recommended BOTS setting remains robust, securing strong TTB and BSF scores, particularly in the aggregated performance metrics. These findings underscore that while domain-specific characteristics may favor different evidence sources, BOTS provides a consistently effective and general-purpose solution across all tested environments.
@@ F.7 Comprehensive Results on Logic
Results on Logic-7B
We extend the evaluation to the Qwen2.5-7B model on the GURU-Logic dataset. Consistent with findings in other settings, BOTS demonstrates robust task selection efficiency. As shown in Figure 19, its ETR curve (red line) shows a rapid initial rise, achieving an effective curriculum faster than other methods. On the downstream benchmarks (Figure 20), the performance of all methods converges, which is an expected outcome when finetuning a powerful base model on these tasks.
This setting, however, reveals two interesting phenomena. First, the curriculum dynamics appear highly sensitive to hyperparameter choices. The ETR trajectories for BOTS and BOTS-MoPPS differ markedly, suggesting a complex interplay between evidence sources in this high-capability, specific-domain context. Second, the offline baseline exhibits a distinct, periodic oscillation in its ETR. A plausible hypothesis for this is that the powerful 7B model rapidly masters clusters of easy tasks within the static curriculum, causing ETR to spike, and then encounters clusters of currently insurmountable tasks, causing ETR to plummet. This cycle repeats as the model iterates over the fixed dataset, resulting in the observed sharp fluctuations.

### v3

@@ G.1 Computational Overhead
G.2 Interpolation-Based Implicit Evidence: Empirical Results
@@
To empirically validate our interpolation-based implicit evidence estimator, we assess its predictive quality against the evolving empirical success probabilities of the training model. We examine the trajectory under vanilla training, i.e., uniformly sampling tasks for training. At each step, we compare the predictions from our interpolation-based estimator with the ground-truth online task success probabilities.
Evaluation Metrics. Two metrics are used: (i) Pearson Correlation, measuring the linear relationship between estimated and empirical difficulties, and (ii) ROC AUC, evaluating the ability to distinguish effective tasks (success strictly between 0 and 1) from ineffective tasks (success equal to 0 or 1).
Training Models. Evaluations are conducted throughout training for two models of different scales: Qwen2.5-1.5B-Instruct and Qwen2.5-7B.
Reference Models. Reference models are selected among three models: Qwen2.5-1.5B-Instruct (1.5B), Qwen2.5-7B-Instruct (7B), and Qwen3-30B-A3B (30B). Besides the default choice of reference models (7B $\times$ 30B) used in Section 4, we additionally include two reference-model pairs: 1.5B $\times$ 7B and 1.5B $\times$ 30B.
The resulting Pearson Correlation and ROC AUC curves are shown in Figure 6.
@@
Figure 6: Performance of the interpolation-based implicit evidence estimator during training of Qwen2.5-1.5B-Instruct (Left) and Qwen2.5-7B (Right), under different choices of reference model pairs: 7B & 30B (Top), 1.5B & 7B (Middle), and 1.5B & 30B (Bottom). Predictive quality is measured by Pearson Correlation and ROC AUC.
Observations.
@@
The consistently positive correlation and ROC AUC above 0.5 demonstrate that interpolation-based implicit evidence effectively captures task difficulty, even in the the extrapolation-heavy setting, where the pair {7B, 30B} is used to predict task success probabilities for Qwen2.5-1.5B-Instruct. All other reference-model combinations maintain even stronger predictive quality throughout training. A mild performance drop is observed in the {1.5B, 30B} $\rightarrow$ 7B setting, which is the only non-extrapolative scenario that both reference models differ substantially in capability from the training model. Additionally, we observe both Pearson Correlation and ROC AUC decline in later stages in several cases, indicating that implicit evidence becomes less informative for difficulty prediction as the model matures. This highlights the necessity of incorporating direct evaluations as explicit evidence to maintain accurate estimation.
Analysis and Takeaways.
These results indicate that choosing well-bracketed reference models—particularly avoiding strong extrapolation regimes and, when possible, including a reference model whose capability is close to the training model—substantially improves the quality of implicit evidence within BOTS.
Figure 7: Qwen2.5-1.5B-Instruct on Math with various combinations of reference models. Ratios of sampled training tasks (measured with 16 rollouts) whose empirical success probabilities are strictly between 0 and 1 (LU), strictly greater than 0 (MU), and strictly less than 1 (RU), together with performance curves averaged over 3 random runs with 95% confidence intervals on MATH500 (LM), AMC23 ($M$), and AIME24 (RM), plotted against training steps. The bottom row shows smoothed versions (running average, window size 3) of the curves in the middle row.
Next, we evaluate how different choices of reference models affect the performance of BOTS. We apply three reference model pairs to BOTS and train Qwen2.5-1.5B-Instruct on the GURU-Math dataset. The results are presented in Figure 7.
Observations.
We observe that the reference model pairs with stronger implicit evidence estimation quality (Figure 6) – namely 1.5B $\times$ 7B and 1.5B $\times$ 30B – achieve clearly better ETR and more efficient training curves compared to the {7B & 30B} pair. Even though 7B $\times$ 30B performs slightly worse than the other combinations, it still delivers performance that significantly surpasses the random baseline.
Analysis and Takeaways.
Overall, BOTS exhibits low sensitivity to the choice of reference models: even imperfect implicit-evidence estimates (e.g., under extrapolation) still yield substantial training improvements. At the same time, better-chosen reference models produce higher-quality implicit evidence, which in turn further strengthens training efficiency and end performance.
G.3 Sampling from Posterior
@@
G.4 Dynamics of Selected Tasks
@@
G.5 External Baseline: Dynamic Sampling
Figure 11: Qwen2.5-1.5B-Instruct on Math. Performance curves averaged over 3 random runs with 95% confidence intervals on MATH500 (LU), AMC23 (MU), and AIME24 (RU), plotted against training steps. The bottom row shows smoothed versions (running average, window size 3) of the curves in the upper row.
[…]
Observations.
Interestingly, while Dynamic Sampling consistently outperforms the random baseline, it slightly underperforms BOTS even under a step-wise comparison—despite achieving the ideal effective task ratio (ETR = 1).
Analysis and Takeaways.
A likely explanation is that Dynamic Sampling only filters out invalid tasks (success probability 0 or 1) but does not differentiate among the remaining ones. In contrast, BOTS explicitly prioritizes tasks whose success probabilities lie near the ideal difficulty (e.g., 0.5), which are most beneficial for learning. Consequently, BOTS achieves both higher step-wise efficiency and lower computational cost. Overall, in the studied setting, BOTS outperforms Dynamic Sampling on both axes.
G.6 Extended Experimental Results: Impact of $\rho$
Figure 12: Qwen2.5-1.5B-Instruct on Math with $\lambda=0.1,\rho\in\{0.0,0.05,0.1,0.2,0.5,1.0\}$ Ratios of sampled training tasks (measured with 16 rollouts) whose empirical success probabilities are strictly between 0 and 1 (LU), strictly greater than 0 (MU), and strictly less than 1 (RU), together with performance curves averaged over 3 random runs with 95% confidence intervals on MATH500 (LM), AMC23 ($M$), and AIME24 (RM), plotted against training steps. The bottom row shows smoothed versions (running average, window size 3) of the curves in the middle row.
G.7 Additional Experimental Results: Impact of $\lambda$
Figure 13: Qwen2.5-1.5B-Instruct on Math with $\lambda\in\{0.0,0.05,0.1,0.2,0.5,1.0\},\rho=0.1$. Ratios of sampled training tasks (measured with 16 rollouts) whose empirical success probabilities are strictly between 0 and 1 (LU), strictly greater than 0 (MU), and strictly less than 1 (RU), together with performance curves averaged over 3 random runs with 95% confidence intervals on MATH500 (LM), AMC23 ($M$), and AIME24 (RM), plotted against training steps. The bottom row shows smoothed versions (running average, window size 3) of the curves in the middle row.
G.8 Extended Experimental Results: Math
Figure 14: Qwen2.5-1.5B-Instruct on Math. Ratios of sampled training tasks (measured with 16 rollouts) whose empirical success probabilities are strictly between 0 and 1 (LU), strictly greater than 0 (MU), and strictly less than 1 (RU), together with performance curves averaged over 3 random runs with 95% confidence intervals on MATH500 (LM), AMC23 ($M$), and AIME24 (RM), plotted against training steps. The bottom row shows smoothed versions (running average, window size 3) of the curves in the middle row.
Benchmark MATH500 AMC23 AIME24 Math Aggregated Performance
Target Fraction 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100% 50% 75% 100% 25% 50% 100%
Offline 0.76 0.67 0.88 1.08 1.10 1.02 0.89 0.65 0.90 1.11 1.13 1.06 0.93 0.77 0.45 1.07 1.26 1.35 0.77 0.66 0.85 1.09 1.11 1.04
BOTS-MoPPS 0.98 0.94 - 0.99 1.00 0.99 1.20 0.85 - 0.98 1.09 0.97 1.26 1.41 0.77 0.95 0.93 1.15 1.02 0.90 0.89 0.98 1.03 1.00
BOTS-DOTS 0.72 0.56 0.74 1.12 1.07 1.01 0.75 0.76 0.66 1.03 1.08 1.01 0.88 1.23 0.64 1.18 0.96 1.24 0.70 0.70 0.73 1.08 1.08 1.01
BOTS 0.86 0.66 0.78 1.05 1.09 1.06 0.86 0.68 0.74 1.10 1.16 1.06 0.86 0.85 0.50 1.20 1.25 1.23 0.85 0.66 0.72 1.06 1.12 1.05
Table 8: TTB and BSF evaluated on Math with Qwen2.5-1.5B-Instruct. For TTB, notation “-” indicates that the target performance is never achieved within the evaluation window. The best and second best results are marked accordingly.
@@ G.8 Extended Experimental Results: Math
G.9 Extended Experimental Results: Code
@@
G.10 Extended Experimental Results: Logic
@@
Appendix H Usage of Large Language Models
We employed Large Language Models solely for the purpose of polishing the writing in this manuscript. Their function was limited to tasks such as correcting grammatical errors, rephrasing sentences to enhance clarity and flow, and ensuring the consistent use of terminology. The LLMs had no role in the ideation of the research, the development of the BOTS framework, the experimental design, or the analysis of results.